OpenStudy (anonymous):

You roll two dice. What is the probability of rolling a 2 or a 4 on the second die, given that you rolled an even number on the first die

OpenStudy (anonymous):

The sample space when you roll two dice: $$ \begin{array}{c|c|c|C|} &1&2&3&4&5&6 \\ \hline 1 &(1,1)&(1,2)&(1,3)&(1,4)&(1,5)&(1,6)\\ \hline 2 &(2,1)&(2,2)&(2,3)&(2,4)&(2,5)&(2,6)\\ \hline 3 &(3,1)&(3,2)&(3,3)&(3,4)&(3,5)&(3,6)\\ \hline 4 &(4,1)&(4,2)&(4,3)&(4,4)&(4,5)&(4,6)\\ \hline 5 &(5,1)&(5,2)&(5,3)&(5,4)&(5,5)&(5,6)\\ \hline 6 &(6,1)&(6,2)&(6,3)&(6,4)&(6,5)&(6,6)\\ \hline \end{array}$$

OpenStudy (anonymous):

Well, first of all, does the result of the first die affect what you'll get on the second die?

OpenStudy (anonymous):

Can you spot your desired ones?

OpenStudy (anonymous):

Check out the column 2 and 4.

OpenStudy (lgbasallote):

that's some heck of latex

OpenStudy (anonymous):

well the full question is this: You roll two dice. What is the probability of rolling a 2 or a 4 on the second die, given that you rolled an even number on the first die? A 6X6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

I just posted that \( 6\times 6\) table for you.

OpenStudy (anonymous):

the thing is i don't know how to figure it out, even with the chart

OpenStudy (anonymous):

help please!

OpenStudy (anonymous):

Do you understand conditional probability?

OpenStudy (anonymous):

i don't understand conditional probability.

OpenStudy (anonymous):

there's 6 pairs? so 6 out of 36? is 1/6?

OpenStudy (anonymous):

Let A is an event where you get an even number on the first die. Let B is the event of rolling 2 or a 4 on the second die. Can you spot the the pairs satisfying the event A and B respectively?

OpenStudy (anonymous):

A=\(\{(2,1),(2,2),(2,3),(2,4),(2,5),(2,6)\cdots (6,6)\} \) There would be 18 such cases. B= all the elements under the column 2 and 4 respectively.

OpenStudy (anonymous):

There will be 12 such cases.

OpenStudy (anonymous):

so 18/36 which is 1/2 ?

OpenStudy (anonymous):

wait 12? i thought you said 18?

OpenStudy (anonymous):

\(A\cap B = ?\)

OpenStudy (anonymous):

ok so there are 2 dice and the possibilites are 36. but out of those only 12 are effective. so 12/36= is 1/3 ?

OpenStudy (kropot72):

The probability of an even number on A is 3/6 and the probability of a 2 or a 4 on B is 2/6. The compound event is made up out of 2 separate and independent events. Therefore the probability of the compound event is the product of the separate probabilities: \[P=\frac{3}{6}\times \frac{2}{6}=\frac{6}{36}\]